OpenStudy (anonymous):

another long one. the speed of train A is 14 mph slower than train B. Train A travels 200 miles the same time that it takes B to travel 270 miles. What are the speeds of each train. Help.

OpenStudy (anonymous):

well,the times are equal, correct? and distance = speed * time. Where do you go from there?

OpenStudy (anonymous):

im honestly lost :(

OpenStudy (anonymous):

I will try to explain this a little better for you if you are having trouble. Let Da = distance traveled by train A Db = distance traveled by train b Va = speed of train A Vb = speed of train B t = time okay so you can start out using distance = speed * time, or d=vt times the trains spent traveling are both equal, so rearranging the equations: Da=Va*t Db=Vb*t into: t=Da/Va t=Db/Vb will allow you to set them equal to each other, giving you: Da/Va=Db/Vb. It was said in the problem that Train a was moving 14 mph slower, so lets change our Va and Vb to reflect this Vb = V Va= (V-14) and substitute your distance values Da = 200 mi Db = 270 mi that gives us: 200/(V-14) = 270/V 200V=270(V-14) 200V = 270V - 3780 70V = 3780 that would mean V = 54mph since V = Vb, Vb = 54 mph. Va = 40 mph

OpenStudy (anonymous):

thanks alot